OpenStudy (anonymous):

Factor the trinomial: x^2 + 22x + 72 A. (x + 2)(x + 11) B. (x + 4)(x + 18) C. (x + 2)(x + 36) D. (x + 6)(x + 12)

OpenStudy (shamil98):

What values in (x+b)(x+b) , add up to 22, and multiply to 72? , What do you think the factored form is?

OpenStudy (anonymous):

go to www.mathway.com and you can type it in and it'll do it for you

OpenStudy (anonymous):

what multiplies to 72 and adds to 22

OpenStudy (anonymous):

1 72 2 36 3 24 4 18 4+18 =22 4*18=72 (x+4)(x+18)